# Diffraction Aided Wireless Positioning

### v1

@@ II-B Path Length Model
In this section, we apply the law of diffraction to derive a closed-form expression for the path length $AQ_{e}B$, which is presented in the next lemma.
@@
The path length $p$ traced out by the geometrical path $AQ_{e}B$ in Fig. 1(a) is expressed as
@@
Here, the coordinates of the diffraction point $Q_{e}$, $\bm{Q}_{e}=[q_{x},0,z_{e}]^{T}$ can be expressed as a convex combination of the coordinates of the endpoints of the edge $\bm{X}_{1}=[x_{1},0,z_{e}]^{T}$ and $\bm{X}_{2}=[x_{2},0,z_{e}]^{T}$ as
@@
Remark.
This result can also be derived from Fermat’s principle of least time. By applying this principle, we calculate the derivative of the path length expression in eq. (3) with respect to the x-coordinate $q_{x}$ of the diffraction point $Q_{e}$ and set the expression obtained to zero i.e., $\frac{\delta p}{\delta q_{x}}=0$. This gives the same expression that is obtained from the diffraction law.
@@ II-C Diffraction Electric Field
We derive an expression for the received diffraction electric field at the observation point $B$ in Fig. 1(a). The electric field polarization along the incident ray is resolved into two perpendicular components $\hat{\beta}_{0}^{\prime}$, $\hat{\phi}^{\prime}$ both of which are also perpendicular to the propagation direction of the incident ray $\hat{s^{\prime}}$. Similarly, the electric field of the diffracted ray is resolved into $\hat{\bm{\beta}}_{0}$, $\hat{\bm{\phi}}$ components, both being perpendicular to the propagation direction of the diffracted ray $\hat{\bm{s}}$. We now define additional vectors to define two new coordinate systems called the ray fixed coordinate system and the edge fixed coordinate system as
$\begin{split}\hat{\phi}^{\prime}=\frac{-\hat{e}\times\hat{s}^{\prime}}{\left|\hat{e}\times\hat{s}^{\prime}\right|},\;\hat{\beta}_{0}^{\prime}=\hat{\phi}^{\prime}\times\hat{s}^{\prime},\;\hat{\phi}=\frac{\hat{e}\times\hat{s}}{|\hat{e}\times\hat{s}|},\;\hat{\beta}_{0}=\hat{\phi}\times\hat{s}.\end{split}$ (5)
@@
The diffracted field $\overrightarrow{E}^{d}$ due to the edge $\hat{e}$ of the half-plane can be expressed along the orthogonal vectors $\hat{\beta}_{0}$ and $\hat{\phi}$ as
$\overrightarrow{E}^{d}=E_{\beta_{0}}^{d}\hat{\beta}_{0}+E_{\hat{\phi}}^{d}\hat{\phi}.$ (6)
Here, $E_{\beta_{0}}^{d}$ and $E_{\phi}^{d}$ are the diffracted electric field components along the $\hat{\beta}_{0}$ and $\hat{\phi}$ directions and are obtained using
$\left[\begin{array}[]{c}{E}_{\beta_{0}}^{d}\\ {E}_{\phi}^{d}\end{array}\right]=\left[\begin{array}[]{cc}-D_{s}&0\\ 0&-D_{h}\end{array}\right]\left[\begin{array}[]{c}{E}_{{\beta_{0}}^{\prime}}^{i}\left(Q_{e}\right)\\ {E}_{\phi^{\prime}}^{i}\left(Q_{e}\right)\end{array}\right]\frac{\mathrm{e}^{-jk|s|}}{\sqrt{|s|}}.$ (7)
Here, $D_{s}$ and $D_{h}$ are the soft and hard diffraction coefficients, respectively, $|s|$ is the distance between the field point $B$ and the diffraction point $Q_{e}$ and $k$ is the wavenumber. ${E}_{{\beta_{0}}^{\prime}}^{i}\left(Q_{e}\right)$ and ${E}_{\phi^{\prime}}^{i}\left(Q_{e}\right)$ are the components of the incident electric field at the point $Q_{e}$ along $\hat{\beta}_{0}^{\prime}$ and $\hat{\phi}^{\prime}$ directions respectively. The incident field components at point $Q_{e}$ are calculated using
$\begin{split}{E}_{{\beta_{0}}^{\prime}}^{i}\left(Q_{e}\right)=\left(\overrightarrow{E}_{0}\cdot\hat{\beta}_{0}^{\prime}\right)\frac{\mathrm{e}^{-jk|s^{\prime}|}}{|s^{\prime}|},\\ {E}_{\phi^{\prime}}^{i}\left(Q_{e}\right)=\left(\overrightarrow{E}_{0}\cdot\hat{\phi}^{\prime}\right)\frac{\mathrm{e}^{-jk|s^{\prime}|}}{|s^{\prime}|}.\end{split}$ (8)
Here, $\overrightarrow{E}_{0}$ is the linearly polarized transmitted electric field polarization vector at the source point $A$, $|s^{\prime}|$ is the distance from the source point $A$ to the diffraction point $Q_{e}$ and $\cdot$’ is the vector dot product. Finally, the soft ($D_{s}$) and hard ($D_{h}$) Keller Diffraction coefficients for a thin sheet are given by [21, 22]
@@
The angles $\psi^{\prime}$ and $\psi$ are obtained using the procedure below. We first define unit vectors $\hat{s}_{t}^{\prime}$ and $\hat{s}_{t}$ as
$\begin{split}\hat{s}_{t}^{\prime}=\frac{\hat{s}^{\prime}-\left(\hat{s}^{\prime}\cdot\hat{e}\right)\hat{e}}{\left|\hat{s}^{\prime}-\left(\hat{s}^{\prime}\cdot\hat{e}\right)\hat{e}\right|},\;\;\hat{s}_{t}=\frac{\hat{s}-(\hat{s}\cdot\hat{e})\hat{e}}{\left|\hat{s}-(\hat{s}\cdot\hat{e})\hat{e}\right|},\end{split}$ (10)
@@
$\begin{split}\psi^{\prime}&=\pi-\left[\pi-\arccos\left(-\hat{s}_{t}^{\prime}\cdot\hat{t}_{o}\right)\right]\operatorname{sgn}\left(-\hat{s}_{t}^{\prime}\cdot\hat{n}_{o}\right),\\ \psi&=\pi-\left[\pi-\arccos\left(\hat{s}_{t}\cdot\hat{t}_{o}\right)\right]\operatorname{sgn}\left(\hat{s}_{t}\cdot\hat{n}_{o}\right).\end{split}$ (11)
Note, $0\leq\arccos{x}\leq\pi$ and the function $\operatorname{sgn}\left(x\right)$ refers to the standard signum function definition. These diffraction coefficients were initially proposed by Keller [20] and are valid for regions away from the shadow boundaries. However, they are discontinuous over the shadow boundaries, and to resolve the discontinuity, they can be replaced with the UTD diffraction coefficients [29]. Since UTD and GTD coefficients agree for regions away from the shadow boundaries [22], this assumption does not impact our results.
@@ III Application to an O2I Scenario
Consider Fig. 1 for an emergency scenario with emergency responders at unknown locations inside a multi-floored building. Given the fact that GPS is not reliable indoors and that in-building WiFi may be inoperabe during an emergency, consider a candidate positioning system consisting of UAVs that are quickly deployed around the affected building. These UAVs form a wireless network, have position knowledge, and behave as anchors at known locations. UAVs offer several advantages including reduced reliance on fixed infrastructure, making the positioning system resilient to damaged infrastructure during emergencies [30]. Further, the mobility of UAVs can be leveraged to improve indoor coverage for communication purposes by placing them in advantageous positions close to the building [19] and serving positioning needs [18]. However, there is still a requirement to improve especially the Z-axis positioning accuracy [17].
Kohli et al. [31] conducted a path-loss based analysis at 28 GHz for the O2I scenario and concluded the O2I path loss depended on the type of glass on the building exterior. Bas et al. [32] conducted a measurement campaign for a similar O2I scenario where they investigated the signal propagation from an outdoor transmitter to an indoor receiver. They concluded that at 28 GHz for brick buildings, the direct path between the transmitter and receiver if blocked by the exterior of the building, is severely attenuated. Further, they concluded that the MPCs observed at indoor locations resulted from interactions with the windows. Our analysis of the diffraction paths in Section II-A based on Fermat’s principle of least time suggests that the first arriving paths at indoor locations are the result of diffraction from the window edges. Further, if we can estimate the path lengths of these diffraction paths, we could circumvent the problem of NLOS bias, thereby improving both 3D and Z-axis position estimates.
Our analysis starts by developing a simplified model of the building to model the diffraction MPCs. Based on this model, we develop insights about the signal propagation from the outdoor anchor to the indoor node. We also present results from realistic EM modeling software that uses raytracing to investigate the different signal propagation mechanisms and validate some of our insights.
@@ III-A Simplified Model of a Building
Our simplified model of the building begins by assuming that every building floor can be modeled as two half-planes separated by a vertical distance $w$ corresponding to the vertical dimension of the window as shown in Fig. 1(b). Consider a node that is located on an arbitrary floor of the building at $\bm{\alpha}=[x_{n},y_{n},z_{n}]^{T}$. The transmit signal originates from an anchor at a known location $\bm{X}_{a}=[x_{a},y_{a},z_{a}]^{T}$ below the building floor where the node is located, such that there is NLOS signal propagation. At the node location, we would obtain two diffraction MPCs resulting from the diffraction from the upper and lower edges of the window located on the same floor of the building. Note that our model does not contain vertical pillars separating windows on the same floor. This is because, on applying Lemma 1 for the vertical edges located on the vertical pillars, we would obtain $\lambda>1$ or $\lambda<0$ on solving the corresponding quadratic equation. This implies that the diffraction point $Q_{e}$ lies beyond the endpoints of the vertical edges. In this case, the diffraction would be from the corner where the window’s horizontal and vertical edges meet. This phenomenon is called corner diffraction and in our work we ignore it since it is known to be significantly weaker as compared to edge diffraction [33].
An important observation is the diffraction MPCs are generated from the edges of the window located on the same floor as the node. This suggests that Z-axis position information is inherently encoded in the diffraction MPCs. Our goal is to obtain the path length and the electric field for these two diffraction MPCs from the upper and lower window edges.We use superscripts ‘$u$’ and ‘$l$’ to refer to the upper and lower diffraction MPC parameters respectively.
@@
In our simplified model of a building shown in Fig. 1(b), the active diffraction edge from which the diffraction MPCs are generated depends on which floor the node is located on. Therefore, the z-coordinate of the node and the diffracting edges are always separated by a constant offset. Without loss of generality, we assume the vertical coordinate of the node $z_{n}$ is located at the midpoint of node building floor. Therefore, $z_{n}-z_{e}^{u}=-0.5w$ for the upper edge diffraction MPC and $z_{n}-z_{e}^{l}=0.5w$ for the lower edge diffraction MPC. Here, $z_{e}^{u}$ and $z_{e}^{l}$ are the vertical coordinates of the upper and lower diffracting edges, respectively.
The distance of the node’s z-coordinate from the diffracting edges depends on the mounting height of the receiver and vertical separation of the diffracting edges and will remain fixed for a given building and mounting height. This assumption is used only to arrive at the analytical results and eventually only leads to a fixed error in the Z-axis estimate. However, this constant offset could be suitably incorporated in a pre-deployment calibration process and only complicates further analysis and thus has been neglected.
@@
Two diffraction MPCs originate from an anchor location $A$ outside the building to a node location $B$ on an arbitrary floor inside the building. These two MPCs are due to diffraction from the lower and upper edge of the window of vertical dimensions $w$ situated on the same floor as the node. For each MPC, assuming the diffraction point $Q_{e}$ lies on the respective diffracting edge, the path length can be expressed as the sum of two Euclidean distances $|AQ_{e}|$ as the Outside Path Length (OPL) and $|Q_{e}B|$ as the Inside Path Length (IPL). The path length expression for the upper edge MPC is
$\begin{split}p(\bm{\alpha})=&\underbrace{\sqrt{(x_{a}-q_{x})^{2}+(y_{a})^{2}+(z_{a}-z_{n}+0.5w)^{2}}}_{OPL}\\ &+\underbrace{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}}_{IPL}.\\ \end{split}$ (12)
Here, the diffraction point coordinates for the upper edge diffraction path are $[q_{x},0,z_{n}+0.5w]^{T}$. Note, $q_{x}$ can be obtained using eq. (4).
@@
We substitute the value of the z-coordinate of the upper edge in Assumption 1 into the path length expression in Lemma 1 to obtain a closed form expression for the path length of the upper edge diffraction MPC. Note that we need to substitute the appropriate coordinates of the endpoints $\bm{X}_{1}^{u}$ and $\bm{X}_{2}^{u}$ of the upper diffracting edge to calculate the corresponding parameter $\lambda$ to obtain the path length. With similar arguments, we can also get the expression for the path length of the diffraction MPC due to the lower diffraction edge. ∎
We now derive the electric field expressions for the two diffraction MPCs. The electric field for each edge can be calculated using the electric field expression for a half-plane presented in Definition 2. For each edge, we obtain expressions for the incident ray vector $\hat{\bm{s}}^{{}^{\prime}}$, diffracted ray vector $\hat{\bm{s}}$, and the edge vector $\hat{\bm{e}}$ and the final expressions are presented in Table I.
TABLE I: Parameters for upper and lower edges for $\hat{\bm{e}}$, $\hat{\bm{s}}^{{}^{\prime}}$ and $\hat{\bm{s}}$.
[…]
Lower edge $q_{x}\hat{x}+0\hat{y}+(z_{e}^{l})\hat{z}$ $-\hat{y}$ $-\hat{z}$ $-\hat{x}$ $\frac{(q_{x}-x_{a})\hat{x}-(y_{a})\hat{y}+(z_{e}^{l}-z_{a})\hat{z}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}}$ $\frac{(x_{n}-q_{x})\hat{x}+(y_{n})\hat{y}+(0.5w)\hat{z}}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}}$
Upper edge $q_{x}\hat{x}+0\hat{y}+(z_{e}^{u})\hat{z}$ $-\hat{y}$ $\hat{z}$ $\hat{x}$ $\frac{(q_{x}-x_{a})\hat{x}-(y_{a})\hat{y}+(z_{e}^{u}-z_{a})\hat{z}}{\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}}$ $\frac{(x_{n}-q_{x})\hat{x}+(y_{n})\hat{y}-(0.5w)\hat{z}}{\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}}$

### v2

@@ II-B Path Length Model
In this section, we apply the law of diffraction to derive a closed-form expression for the path length $\bm{X_{a}Q_{e}\alpha}$, which is presented in the next lemma.
@@
The path length $p$ traced out by the geometrical path $\bm{X_{a}Q_{e}\alpha}$ in Fig. 1(a) is expressed as
@@
Here, the coordinates of the diffraction point $\bm{Q}_{e}=[q_{x},0,z_{e}]^{T}$, can be expressed as a convex combination of the coordinates of the endpoints of the edge $\bm{X}_{1}=[x_{1},0,z_{e}]^{T}$ and $\bm{X}_{2}=[x_{2},0,z_{e}]^{T}$ as
@@
This result can also be derived from Fermat’s principle of least time. By applying this principle, we calculate the derivative of the path length expression in eq. (3) with respect to the x-coordinate $q_{x}$ of the diffraction point $\bm{Q_{e}}=[q_{x},0,z_{e}^{u}]^{T}$ and set the expression obtained to zero, that is, $\frac{\delta p}{\delta q_{x}}=0$. This gives the same expression as is obtained from the diffraction law.
This is a general path length expression applicable for modeling diffraction from any edge. In the conference version of this article [1], we used this expression for indoor positioning; however, it required detailed positional knowledge of the diffracting edges. Corollary 1 in section III-A presents an improved diffraction path length expression tailored to the modeled scenario, and in Corollary 3 in section IV-B we derive the necessary and sufficient conditions for positioning using this model. Thus, we are able to eliminate the need for detailed positioning knowledge about the diffraction edges.
@@ II-C Diffraction Electric Field
We derive an expression for the received diffraction electric field at the observation point $\bm{\alpha}$ in Fig. 1(a). The electric field polarization along the incident ray is resolved into two perpendicular components $\bm{\hat{\beta}_{0}^{\prime}}$, ϕ^superscriptbold-^bold-italic-$\bm{\hat{\phi}^{\prime}}$ both of which are also perpendicular to the propagation direction of the incident ray $\bm{\hat{s^{\prime}}}$. Similarly, the electric field of the diffracted ray is resolved into $\bm{\hat{\bm{\beta}}_{0}}$, $\bm{\hat{\phi}}$ components, both perpendicular to the propagation direction of the diffracted ray $\bm{\hat{s}}$. We now define additional vectors to define two new coordinate systems, called the ray fixed coordinate system and the edge fixed coordinate system.
$\begin{split}\bm{\hat{\phi}^{\prime}}=\frac{-\bm{\hat{e}}\times\bm{\hat{s}^{\prime}}}{\left|\bm{\hat{e}}\times\bm{\hat{s}^{\prime}}\right|},\;\bm{\hat{\beta}_{0}^{\prime}}=\bm{\hat{\phi}^{\prime}}\times\bm{\hat{s}^{\prime}},\;\bm{\hat{\phi}}=\frac{\bm{\hat{e}}\times\bm{\hat{s}}}{|\bm{\hat{e}}\times\bm{\hat{s}}|},\;\bm{\hat{\beta}_{0}}=\bm{\hat{\phi}}\times\bm{\hat{s}}.\end{split}$ (5)
@@
The diffracted field $\bm{\overrightarrow{E}^{d}}$ due to the edge $\bm{\hat{e}}$ of the half-plane can be expressed along two orthogonal vectors $\bm{\hat{\beta}_{0}}$ and $\bm{\hat{\phi}}$ as
$\bm{\overrightarrow{E}^{d}}=E_{\beta_{0}}^{d}\bm{\hat{\beta}_{0}}+E_{\hat{\phi}}^{d}\bm{\hat{\phi}}.$ (6)
Here, $E_{\beta_{0}}^{d}$ and $E_{\phi}^{d}$ are the diffracted electric field components along the $\bm{\hat{\beta}_{0}}$ and $\bm{\hat{\phi}}$ directions and are obtained using
$\left[\begin{array}[]{c}{E}_{\beta_{0}}^{d}\\ {E}_{\phi}^{d}\end{array}\right]=\left[\begin{array}[]{cc}-D_{s}&0\\ 0&-D_{h}\end{array}\right]\left[\begin{array}[]{c}{E}_{{\beta_{0}}^{\prime}}^{i}\left(\bm{Q_{e}}\right)\\ {E}_{\phi^{\prime}}^{i}\left(\bm{Q_{e}}\right)\end{array}\right]\frac{\mathrm{e}^{-jk|\bm{s}|}}{\sqrt{|\bm{s}|}}.$ (7)
Here, $D_{s}$ and $D_{h}$ are the soft and hard diffraction coefficients, respectively, $|\bm{s}|$ is the distance between the field point $\bm{\alpha}$ and the diffraction point $\bm{Q_{e}}$ and $k$ is the wavenumber. ${E}_{{\beta_{0}}^{\prime}}^{i}\left(\bm{Q_{e}}\right)$ and ${E}_{\phi^{\prime}}^{i}\left(\bm{Q_{e}}\right)$ are the components of the incident electric field at the point $\bm{Q_{e}}$ along $\bm{\hat{\beta}_{0}^{\prime}}$ and ϕ^superscriptbold-^bold-italic-$\bm{\hat{\phi}^{\prime}}$ directions respectively. The incident field components at point $\bm{Q_{e}}$ are calculated using
$\begin{split}{E}_{{\beta_{0}}^{\prime}}^{i}\left(\bm{Q_{e}}\right)=\left(\bm{\overrightarrow{E}_{0}}\cdot\bm{\hat{\beta}_{0}^{\prime}}\right)\frac{\mathrm{e}^{-jk|\bm{s^{\prime}}|}}{|\bm{s^{\prime}}|},\\ {E}_{\phi^{\prime}}^{i}\left(\bm{Q_{e}}\right)=\left(\bm{\overrightarrow{E}_{0}}\cdot\bm{\hat{\phi}^{\prime}}\right)\frac{\mathrm{e}^{-jk|\bm{s^{\prime}}|}}{|\bm{s^{\prime}}|}.\end{split}$ (8)
Here, $\bm{\overrightarrow{E}_{0}}$ is the linearly polarized transmitted electric field polarization vector at the source point $\bm{X_{a}}$, $\bm{|s^{\prime}}|$ is the distance from the source point $\bm{X_{a}}$ to the diffraction point $\bm{Q_{e}}$. Finally, the soft ($D_{s}$) and hard ($D_{h}$) Keller Diffraction coefficients for a thin sheet are given by [26, 25]
@@
The angles $\psi^{\prime}$ and $\psi$ are obtained using the procedure below. We first define unit vectors $\bm{\hat{s}_{t}^{\prime}}$ and $\bm{\hat{s}_{t}}$ as
$\begin{split}\bm{\hat{s}_{t}^{\prime}}=\frac{\bm{\hat{s}^{\prime}}-\left(\bm{\hat{s}^{\prime}}\cdot\bm{\hat{e}}\right)\bm{\hat{e}}}{\left|\bm{\hat{s}^{\prime}}-\left(\bm{\hat{s}^{\prime}}\cdot\bm{\hat{e}}\right)\bm{\hat{e}}\right|},\;\;\bm{\hat{s}_{t}}=\frac{\bm{\hat{s}}-(\bm{\hat{s}}\cdot\bm{\hat{e}})\bm{\hat{e}}}{\left|\bm{\hat{s}}-(\bm{\hat{s}}\cdot\bm{\hat{e}})\bm{\hat{e}}\right|},\end{split}$ (10)
@@
$\begin{split}\psi^{\prime}&=\pi-\left[\pi-\arccos\left(-\bm{\hat{s}_{t}^{\prime}}\cdot\bm{\hat{t}_{o}}\right)\right]\operatorname{sgn}\left(-\bm{\hat{s}_{t}^{\prime}}\cdot\bm{\hat{n}_{o}}\right),\\ \psi&=\pi-\left[\pi-\arccos\left(\bm{\hat{s}_{t}}\cdot\bm{\hat{t}_{o}}\right)\right]\operatorname{sgn}\left(\bm{\hat{s}_{t}}\cdot\bm{\hat{n}_{o}}\right).\end{split}$ (11)
Note, $0\leq\arccos{x}\leq\pi$ and the function $\operatorname{sgn}\left(x\right)$ refers to the standard signum function definition. These diffraction coefficients were initially proposed by Keller [32] and are valid for regions away from the shadow boundaries. However, they are discontinuous over the shadow boundaries and, to resolve the discontinuity, they can be replaced with the UTD diffraction coefficients [36]. Since the UTD and GTD coefficients agree for regions away from the shadow boundaries [25], this assumption does not affect our results.
@@ III Application to an O2I Scenario
Consider Fig. 2 for an emergency scenario with emergency responders at unknown locations inside a multi-floored building. We have a positioning system consisting of UAVs that are quickly deployed around the affected building. These UAVs form a wireless network, have position knowledge, and behave as anchors at known locations. This represents an O2I signal propagation scenario. Kohli et al. [37] conducted a path loss-based analysis at 28 GHz for the O2I scenario and concluded that the O2I path loss depended on the type of glass on the exterior of the building. Bas et al. [38] conducted a measurement campaign for a similar O2I scenario in which they investigated the propagation of the signal from an outdoor transmitter to an indoor receiver. They concluded that at 28 GHz for brick buildings, the direct path between the transmitter and receiver if blocked by the exterior of the building is severely attenuated. In addition, they concluded that the MPCs observed in indoor locations were the result of interactions with windows. Our analysis of the diffraction paths in Section II-A based on Fermat’s principle of least time suggests that the first arriving paths at indoor locations are the result of diffraction from the window edges. Furthermore, if we can estimate the path lengths of these diffraction paths, we could circumvent the problem of NLoS bias, thereby improving both the 3D and z-axis position estimates.
Our analysis starts by developing a simplified model of the building to model the diffraction MPCs. Based on this model, we develop insight into the propagation of the signal from the outdoor anchor to the indoor node. We also present results from realistic electromagnetic modeling software that uses raytracing to investigate different signal propagation mechanisms and validate some of our insights.
@@ III-A Simplified Model of a Building
Our simplified model of the building begins by assuming that every floor of the building can be modeled as two half-planes separated by a vertical distance $w$ corresponding to the vertical dimension of the window, as shown in Fig.,1(b). Consider a node that is located on an arbitrary floor of the building at $\bm{\alpha}=[x_{n},y_{n},z_{n}]^{T}$. The transmit signal originates from the $j^{\text{th}}$ anchor at a known location $\bm{X}_{a,j}=[x_{a,j},y_{a,j},z_{a,j}]^{T}$ below the building floor where the node is located, such that there is NLoS signal propagation. At the node location, we would obtain two diffraction MPCs resulting from the diffraction from the upper and lower edges of the window located on the same floor of the building. Note that our model does not contain vertical pillars that separate windows on the same floor. This is because, applying the lemma 1 for the vertical edges located on the vertical pillars, we would obtain $\lambda>1$ or $\lambda<0$ when solving the corresponding quadratic equation. This implies that the diffraction point $\bm{Q_{e}}$ lies beyond the endpoints of the vertical edges. In this case, the diffraction would be from the corner where the window’s horizontal and vertical edges meet. This phenomenon is called corner diffraction. In our work, we ignore it because it is known to be significantly weaker compared to edge diffraction [39].
An important observation is that the diffraction MPCs are generated from the edges of the window located on the same floor as the node. This suggests that z-axis position information is inherently encoded in the diffraction MPCs. Our goal is to obtain the path length and the electric field for these two diffraction MPCs from the upper and lower window edges. We use superscripts ‘$u$’ and ‘$l$’ to refer to the upper and lower diffraction MPC parameters, respectively.
@@
In our simplified model of a building shown in Fig. 1(b), the active diffraction edges are located on the same building floor as the node. We assume that the z-coordinate of the node and the diffracting edges are always separated by a constant offset. Without loss of generality, we assume that the vertical coordinate of the node $z_{n}$ is located at the midpoint of the node building floor and that the window center coincides with the midpoint of the node building floor. Therefore, $z_{n}=z_{e}^{u}-0.5w$ for the upper edge diffraction MPC and $z_{n}=z_{e}^{l}+0.5w$ for the lower edge diffraction MPC. Here, $z_{e}^{u}$ and $z_{e}^{l}$ are the vertical coordinates of the upper and lower diffracting edges, respectively and $w$ is the vertical dimension of the window.
The distance between the node’s z-coordinate and the diffracting edges generally depends on the receiver’s mounting height above the building floor, the vertical separation of the diffracting edges, and their relative positions with respect to the node. This assumption is used only to arrive at the analytical results. We discuss the impact of relaxing this assumption (D-NLS(Case-1) vs D-NLS (Case-2)) for our proposed positioning technique in the numerical results in Section V-A. This is done by introducing a model mismatch using a vertical offset $\Delta$ in eq. (23).
@@
Two diffraction MPCs originate from an anchor location $\bm{X_{a,j}}$ outside the building to a node location $\bm{\alpha}$ on an arbitrary floor inside the building. These two MPCs are due to diffraction from the lower and upper edges of the window with vertical dimensions $w$. This window is situated on the same building floor as the node. For each MPC, assuming that the diffraction point $\bm{Q_{e}}$ lies on the respective diffracting edge, the path length can be expressed as the sum of two Euclidean distances $|\bm{X_{a,j}Q_{e}}|$ as the Outside Path Length (OPL) and $|\bm{Q_{e}\alpha}|$ as the Inside Path Length (IPL). The path length expression for the upper edge MPC is
$\begin{split}p_{j}(\bm{\alpha})=&\underbrace{\sqrt{(x_{a,j}-q_{x,j}^{u})^{2}+(y_{a,j})^{2}+(z_{a,j}-z_{n}-0.5w)^{2}}}_{OPL}\\ &+\underbrace{\sqrt{(x_{n}-q_{x,j}^{u})^{2}+(y_{n})^{2}+(0.5w)^{2}}}_{IPL}.\\ \end{split}$ (12)
Here, the diffraction point coordinates for the upper edge diffraction path are $\bm{Q_{e}^{u}}=[q_{x}^{u},0,z_{n}+0.5w]^{T}$. Note that $q_{x}^{u}$ can be obtained using eq. (4).
@@
We substitute the value of the z-coordinate of the upper edge in the Assumption 1 into the expression of the path length in Lemma 1 to obtain a closed-form expression for the path length of the upper edge diffraction MPC. Note that we need to substitute the appropriate coordinates of the end points $\bm{X}_{1}^{u}$ and $\bm{X}_{2}^{u}$ of the upper diffracting edge to calculate the corresponding parameter $\lambda$ to obtain the path length. With similar arguments, we can also get the expression for the path length of the diffraction MPC due to the lower diffraction edge. ∎
We now derive the electric field expressions for the two diffraction MPCs. The electric field for each edge can be calculated using the electric field expression for a half-plane presented in Definition 2. For each edge, we obtain expressions for the incident ray vector $\bm{\hat{s}^{{}^{\prime}}}$, diffracted ray vector $\bm{\hat{s}}$, and the edge vector $\bm{\hat{e}}$ and the final expressions are presented in Table I.
TABLE I: Parameters for upper and lower edges for $\bm{\hat{e}}$, $\bm{\hat{s}^{{}^{\prime}}}$ and $\bm{\hat{s}}$.
$\bm{Q_{e}}$ $\bm{\hat{n}_{0}}$ $\bm{\hat{t}_{0}}$ $\bm{\hat{e}}$ $\bm{\hat{s}^{\prime}}$ $\bm{\hat{s}}$
Lower edge $q_{x}^{l}\bm{\hat{x}}+0\bm{\hat{y}}+(z_{e}^{l})\bm{\hat{z}}$ $-\bm{\hat{y}}$ $-\bm{\hat{z}}$ $-\bm{\hat{x}}$ $\frac{(q_{x}^{l}-x_{a})\bm{\hat{x}}-(y_{a})\bm{\hat{y}}+(z_{e}^{l}-z_{a})\bm{\hat{z}}}{\sqrt{(q_{x}^{l}-x_{a})^{2}+(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}}$ $\frac{(x_{n}-q_{x}^{l})\bm{\hat{x}}+(y_{n})\bm{\hat{y}}+(0.5w)\bm{\hat{z}}}{\sqrt{(x_{n}-q_{x}^{l})^{2}+(y_{n})^{2}+(0.5w)^{2}}}$
Upper edge $q_{x}^{u}\bm{\hat{x}}+0\bm{\hat{y}}+(z_{e}^{u})\bm{\hat{z}}$ $-\bm{\hat{y}}$ $\bm{\hat{z}}$ $\bm{\hat{x}}$ $\frac{(q_{x}^{u}-x_{a})\bm{\hat{x}}-(y_{a})\bm{\hat{y}}+(z_{e}^{u}-z_{a})\bm{\hat{z}}}{\sqrt{(q_{x}^{u}-x_{a})^{2}+(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}}$ $\frac{(x_{n}-q_{x}^{u})\bm{\hat{x}}+(y_{n})\bm{\hat{y}}-(0.5w)\bm{\hat{z}}}{\sqrt{(x_{n}-q_{x}^{u})^{2}+(y_{n})^{2}+(0.5w)^{2}}}$